# Text-to-Decision Agent: Learning Generalist Policies from Natural Language Supervision

### v2

@@ Appendix F Analysis of Parameter-Efficient Fine-tuning
Appendix G More Visualization Insights
We also visualize text and decision embeddings of the Cheetah-Vel environment through t-SNE, as shown in Figure 8.
Figure 8: t-SNE visualization of Cheetah-Vel where tasks with target velocities in $[0.075,3.0]$ are mapped into rainbow-colored points. Top: the evolution from Raw Decision Embeddings to Dynamics-Aware Decision Embeddings, where the initially entangled trajectories are transformed into well-separated clusters that highlight the successful capture of task-specific environment dynamics. Bottom: using dynamics-aware decision embeddings to align text embeddings with contrastive loss. Raw Text Embeddings are scattered with a vague distribution structure. In contrast, Aligned Text Embeddings form a clear rectilinear distribution from red (low velocity) to blue (high velocity), exactly matching the rectilinear spectrum of target velocities in a physical sense. This inspiring finding verifies the effective alignment of text embeddings to comprehend environment dynamics and interpretable language grounding in decision domains.

### v3

@@ Appendix E Analysis of Parameter-Efficient Fine-tuning
Appendix F Analysis of Training Pipline
T2DA separately trains three key components: (a) decision embedding, (b) knowledge alignment, and (c) generalist policy. To demonstrate the effectiveness of this modular approach, we conducted ablation studies comparing our method against two alternatives: jointly training (a)+(b), and jointly training all three components (a)+(b)+(c) simultaneously. As shown in Table 9, our separate training pipeline achieves substantial performance gains of 20%-56% across different environments. This modular architecture not only provides a more stable training process but also offers superior scalability, as individual components can be easily substituted or upgraded independently without affecting the entire system.
Table 9: Ablation study results on modular architecture. Best results in bold.
Environment T2DA (a)+(b) (a)+(b)+(c)
Point-Robot $\mathbf{-7.2}$ $-8.9$ $-16.3$
Meta-World $\mathbf{1274.5}$ $1008$ $817.3$
Appendix G Experimental Results on Humanoid
This section presents experimental results on stimulated humanoid robot benchmark[58]. Table 10 shows the converged performance of T2DA and baselines using Mixed datasets under aligned zero-shot setting. The results demonstrate that T2DA consistently outperforms baselines in this more challenging environment.
Table 10: Performance comparison of different methods on Humanoid. Best results in bold and second best underlined.
Method T2DA-T T2DA-D BC-Z DT DD CSRO UNICORN AD
Returns $\mathbf{610.4}$ $\underline{575.4}$ $531.7$ $520.3$ $381.3$ $520.3$ $505.6$ $427.4$